OpenStudy (anonymous):

evaluate -b2 for b=3 b=3, -b2=

OpenStudy (anonymous):

-9 -3^2 = -3 x 3 = -9

OpenStudy (anonymous):

meverett04 is right, but just to clarify: there's a difference between "negative 3 squared" and "the negative of 3 squared." Squaring a number always produces a positive, so (-3)^2 = 9, whereas -(3^2) = -9.

OpenStudy (anonymous):

I see, thanks